A particle moves under the influence of a central force given by F(r) = - k/ rn. If the particle's orbit is circular and passes through the force center, show that n = 5.


1
Expert's answer
2021-12-03T12:44:35-0500

"F(r)=-\frac k{r^n},"

"x^2+y^2=r^2,"

"U=\int Fdr=\frac{k}{(n-1)r^{n-1}},"

"F(r)=-\frac k{r^5}."